# Into the Rabbit Hull: From Task-Relevant Concepts in DINO to Minkowski Geometry

### v1

@@ 7.2 Empirical evidence
We now provide preliminary support for each criterion. The results below should be read as compatible evidence, not proof: multiple mechanisms can mimic the same surface phenomena. Unless stated otherwise we use ImageNet-1k validation tokens from the last DINOv2-B layer; cosine distance and $k$-NN graphs with standard symmetrization. Figure 17 (left) contrasts straight-line interpolation between tokens with piecewise-linear geodesics computed as shortest paths on the token $k$-NN graph. Straight lines depart the data support quickly, whereas graph geodesics remain close throughout. This matches what a Minkowski sum structure predicts: feasible displacements arise as sums of small face-walks within head polytopes, which are piecewise linear in barycentric coordinates yet appear curved in the ambient space. The curved, on-support geodesics thus support criterion (i) by reflecting head-wise convex reweighting rather than simple linear combinations. To test the convex coding assumption (ii), we compare Archetypal Analysis Cutler & Breiman (1994) (AA) with SAE for token reconstruction in Figure˜17 (middle). Note that AA is precisely the single-tile case of MRH ($|S|=1$), making this a direct test of our geometric assumptions. AA imposes dramatically stronger constraints [^9] Archetypal Analysis seeks $\bm{Z}=\bm{X}\bm{B}$ where archetypes are convex combinations of data (so $\bm{Z}\subset\operatorname{conv}(\bm{X})$), and $\bm{X}\approx\bm{Z}\bm{A}$ where data are convex combinations of archetypes. Both $\bm{A}$ and $\bm{B}$ have simplex constraints (columns sum to 1, non-negative).: it forces all reconstructions to lie within the convex hull of observed tokens and requires archetypes to be actual data combinations. Despite these restrictive assumptions, AA matches SAE performance with remarkably few archetypes (10 archetypes per image), providing preliminary evidence that even the simplest case of MRH captures fundamental geometric structure. This is particularly striking in high dimensions, where the probability that a random point lies in the convex hull of a small set of samples decays exponentially with dimension Bárány & Füredi (1988); Balestriero et al. (2021); the fact that tokens can nevertheless be accurately reconstructed from only $\sim$10 archetypes indicates that embeddings concentrate on low-dimensional polytopes embedded within the ambient space. Additionally, the block structure emerges naturally: Figure˜17 (right) reveals that archetypal decompositions spontaneously organize into the block-sparse pattern predicted by criterion (iii), with distinct clusters of co-activating archetypes rather than uniform mixing.
@@ Appendix B From LRH to Dictionary Learning
Here $\mathbf{I}$ is the identity, $\mathbf{e}_{j}$ denotes a canonical basis vector, and $\Psi_{\theta}$ is an encoder and a sparsity projection (e.g., TopK, Jump-ReLU, or simply ReLU) producing sparse codes. This formulation unifies the previous clustering-based concept extraction Ghorbani et al. (2019), orthogonal factorization (PCA/ICE) Zhang et al. (2021), Nonnegative concept extraction (CRAFT/NMF) Fel et al. (2023c), Archetypal Analysis Cutler & Breiman (1994); Thurau & Bauckhage (2009) and modern SAEs Cunningham et al. (2023); Bricken et al. (2023). In practice, these approaches trade off fidelity ($\|\bm{A}-\bm{Z}\bm{D}\|_{F}$) against sparsity (e.g., $\|\bm{Z}\|_{0}$), yielding a Pareto frontier; SAEs are attractive at scale because the encoder $\Psi_{\theta}$ enables amortized, batched inference while retaining LRH’s sparse, overcomplete structure.
@@ Appendix F Baseline Models for Concept Geometry Analysis
Next, to stress test the LRH, we numerically computed the Grassmannian from of $c$ atoms in $d$ dimension. This is a non-trivial computational problem, where analytical solution is only available for few scenarios, otherwise we need to rely on iterative optimization to find approximate Grassmannian frames. Brute force gradient optimization of incoherence is slow to converge at our scale. Here, we adapted the algorithm from recent work TAAP Massion & Massart (2025), to further accelerate the solver, we adapted it to CUDA, made Grassmannian frame solving feasible at our problem scale $c=32000$, $d=768$. To note, even with GPU acceleration, solving the frame once still takes 6hr on an A100 GPU. In the end, we reached maximal coherence of $0.065897$. Indeed, more isotropic and less coherence than the Gaussian random vectors.
@@ I.2 Structure Persists After Position Removal
We have formally defined MRH motivated by an empirical observation of interpolative geometry and now ask a basic question: can a single attention block generate such structure? The answer is yes, and the mechanism is elementary. We first review relevant prior work, then show how attention naturally generates such structure. Previous work provides two lines of supporting evidence. First, neural networks naturally partition input spaces into convex regions through their piecewise-linear activations Montúfar et al. (2014); Raghu et al. (2017); Balestriero et al. (2018), suggesting that convex decompositions may be fundamental to deep architectures Tvetkova et al. (2025); Tankala et al. (2023); Hindupur et al. (2025). Analyses have also approached neural networks explicitly through the polytope lens, showing how piecewise-linear partitions structure representation space Black et al. (2022). Second, recent work on neural population geometry shows that deep networks organize representations into low-dimensional manifolds with rich geometric structure Chung (2021); Cohen et al. (2020), which is consistent with a small number of convex factors combining to yield the observed variability. Third, in the language domain, categorical and hierarchical concepts have been shown to admit convex (polytopal) representations, with semantic relations reflected directly in geometric structure Park et al. (2025).

### v2

@@ 7.2 Empirical evidence
We now provide preliminary support for each criterion. The results below should be read as compatible evidence, not proof: multiple mechanisms can mimic the same surface phenomena. Unless stated otherwise we use ImageNet-1k validation tokens from the last DINOv2-B layer; cosine distance and $k$-NN graphs with standard symmetrization. Figure 17 (left) contrasts straight-line interpolation between tokens with piecewise-linear geodesics computed as shortest paths on the token $k$-NN graph. Straight lines depart the data support quickly, whereas graph geodesics remain close throughout. This matches what a Minkowski sum structure predicts: feasible displacements arise as sums of small face-walks within head polytopes, which are piecewise linear in barycentric coordinates yet appear curved in the ambient space. The curved, on-support geodesics thus support criterion (i) by reflecting head-wise convex reweighting rather than simple linear combinations. To test the convex coding assumption (ii), we compare Archetypal Analysis Cutler and Breiman (1994) (AA) with SAE for token reconstruction in Figure˜17 (middle). Note that AA is precisely the single-tile case of MRH ($|S|=1$), making this a direct test of our geometric assumptions. AA imposes dramatically stronger constraints [^9] Archetypal Analysis seeks $\bm{Z}=\bm{X}\bm{B}$ where archetypes are convex combinations of data (so $\bm{Z}\subset\operatorname{conv}(\bm{X})$), and $\bm{X}\approx\bm{Z}\bm{A}$ where data are convex combinations of archetypes. Both $\bm{A}$ and $\bm{B}$ have simplex constraints (columns sum to 1, non-negative).: it forces all reconstructions to lie within the convex hull of observed tokens and requires archetypes to be actual data combinations. Despite these restrictive assumptions, AA matches SAE performance with remarkably few archetypes (10 archetypes per image), providing preliminary evidence that even the simplest case of MRH captures fundamental geometric structure. This is particularly striking in high dimensions, where the probability that a random point lies in the convex hull of a small set of samples decays exponentially with dimension Bárány and Füredi (1988); Balestriero et al. (2021); the fact that tokens can nevertheless be accurately reconstructed from only $\sim$10 archetypes indicates that embeddings concentrate on low-dimensional polytopes embedded within the ambient space. Additionally, the block structure emerges naturally: Figure˜17 (right) reveals that archetypal decompositions spontaneously organize into the block-sparse pattern predicted by criterion (iii), with distinct clusters of co-activating archetypes rather than uniform mixing.
@@ Appendix B From LRH to Dictionary Learning
Here $\mathbf{I}$ is the identity, $\mathbf{e}_{j}$ denotes a canonical basis vector, and $\Psi_{\theta}$ is an encoder and a sparsity projection (e.g., TopK, Jump-ReLU, or simply ReLU) producing sparse codes. This formulation unifies the previous clustering-based concept extraction Ghorbani et al. (2019), orthogonal factorization (PCA/ICE) Zhang et al. (2021), Nonnegative concept extraction (CRAFT/NMF) Fel et al. (2023c), Archetypal Analysis Cutler and Breiman (1994); Thurau and Bauckhage (2009) and modern SAEs Cunningham et al. (2023); Bricken et al. (2023). In practice, these approaches trade off fidelity ($\|\bm{A}-\bm{Z}\bm{D}\|_{F}$) against sparsity (e.g., $\|\bm{Z}\|_{0}$), yielding a Pareto frontier; SAEs are attractive at scale because the encoder $\Psi_{\theta}$ enables amortized, batched inference while retaining LRH’s sparse, overcomplete structure.
@@ Appendix F Baseline Models for Concept Geometry Analysis
Next, to stress test the LRH, we numerically computed the Grassmannian from of $c$ atoms in $d$ dimension. This is a non-trivial computational problem, where analytical solution is only available for few scenarios, otherwise we need to rely on iterative optimization to find approximate Grassmannian frames. Brute force gradient optimization of incoherence is slow to converge at our scale. Here, we adapted the algorithm from recent work TAAP Massion and Massart (2025), to further accelerate the solver, we adapted it to CUDA, made Grassmannian frame solving feasible at our problem scale $c=32000$, $d=768$. To note, even with GPU acceleration, solving the frame once still takes 6hr on an A100 GPU. In the end, we reached maximal coherence of $0.065897$. Indeed, more isotropic and less coherence than the Gaussian random vectors.
@@ I.2 Structure Persists After Position Removal
We have formally defined MRH motivated by an empirical observation of interpolative geometry and now ask a basic question: can a single attention block generate such structure? The answer is yes, and the mechanism is elementary. We first review relevant prior work, then show how attention naturally generates such structure. Previous work provides two lines of supporting evidence. First, neural networks naturally partition input spaces into convex regions through their piecewise-linear activations Montúfar et al. (2014); Raghu et al. (2017); Balestriero and others (2018), suggesting that convex decompositions may be fundamental to deep architectures Tvetkova et al. (2025); Tankala et al. (2023); Hindupur et al. (2025). Analyses have also approached neural networks explicitly through the polytope lens, showing how piecewise-linear partitions structure representation space Black et al. (2022). Second, recent work on neural population geometry shows that deep networks organize representations into low-dimensional manifolds with rich geometric structure Chung (2021); Cohen et al. (2020), which is consistent with a small number of convex factors combining to yield the observed variability. Third, in the language domain, categorical and hierarchical concepts have been shown to admit convex (polytopal) representations, with semantic relations reflected directly in geometric structure Park et al. (2025).